OpenStudy (anonymous):

there is a building with a 12 ft high window, you want to use a ladder to go up to the window, and you decide to keep the ladder 5 ft away from the building to have a good slant. how long should the ladder be

OpenStudy (anonymous):

...You should be able to attempt this.

OpenStudy (anonymous):

Draw out a diagram, it'll help

OpenStudy (anonymous):



OpenStudy (anonymous):

they really stretch for these word problems don't they imagine a situation where you decide where to put the ladder and then pick from an infinite number of ladders to fit your needs, rather than the other way around yeah, right

OpenStudy (anonymous):

Pythagorean Theorem.

Parth (parthkohli):

What a silly way to write, “Solve \(5^2 + 12^2 = x^2\)

OpenStudy (anonymous):

i cant get the right answer

OpenStudy (blockingthesky):

Well \[5^2 + 12^2 = x^2\] We simplify and get this: \[25 + 144 = x^2\] \[25 + 144 = 169\] and then \[\sqrt{169} = 13 = x\] the answer is 13.

OpenStudy (anonymous):

ty